OpenStudy (anonymous):

What is mean, median, and average of a number? Your help is appreciated. Thank you. With explanations please.

OpenStudy (anonymous):

the mean is when youadd all the numbers together and subtracted by how ever many numbers they have. the median is the number in the middle after you put it in order. and the average is all the numbers added up

OpenStudy (anonymous):

so mean of (12, 10, 5) is =27-3 ?

OpenStudy (anonymous):

Im sorry m,ean is all the numbers added divided by the numbers . let say 12,10,5=27 /3

OpenStudy (anonymous):

ok and whats the difference from average?

OpenStudy (anonymous):

Isnt average alll numbers added and then divided by that many numbers?

OpenStudy (anonymous):

I believe average and mean CAN be the same thing from my experience also avg = (1+2+3)/n

OpenStudy (anonymous):

thanks guys:)

OpenStudy (anonymous):

mean is average median is the middle number